deposits are made into an account that earns 7% compounded annually. deposits of $100 each were made on March 1,2015 and September 1, 2016. if interest is calculated on January 1st determine the balance of the account on January 1, 2018
Answer #1

From March 1 2015 to January 1 2018 number of Month = 2 year 10 months = 34/12 years
From Sep 1 2016  to January 1 2018 number of Month = 1 year 4 months = 4/3 years

The Balance of account in Jan 1, 2018 = 100*(1+7%)34/12 + 100*(1+7%)4/3 = 230.57

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well